You have 20% saline and 50% saline, how much of each solution should be mixed together to get 60mL of 28% saline?

Answer by mananth(16946)

percent ---------------- quantity
Saline I 20.00% ---------------- x ml
saline II 50.00% ------ 60 - x ml
Mixture 28.00% ---------------- 60
Total 60 ml
20.00% x + 50.00% ( 60 - x ) = 28.00% * 60
20 x + 50 ( 60 - x ) = 1680
20 x + 3000 - 50 x = 1680
20 x - 50 x = 1680 - -3000
-30 x = -1320
/ -30
x = 44 ml 20.00% Saline I
16 ml 50.00% saline II

m.ananth@hotmail.ca